OpenStudy (anonymous):

The equation y = 11x shows that a boat travels at a speed of 11 km/h. Which ordered pair satisfies the equation y = 11x? A. (0, 11) B. (1, 12) C. (2, 22) D. (11, 1)

OpenStudy (freckles):

what point can you plug in and the equation hold? for example which equation is true: 11=11(0) 12=11(1) 22=11(2) 1=11(11)

OpenStudy (anonymous):

1

OpenStudy (freckles):

which equation are you picking?

OpenStudy (freckles):

11 times 0=? 11 times 1=? 11 times 2=? 11 times 11=?

OpenStudy (anonymous):

11 * 1

OpenStudy (anonymous):

Plug the ordered pairs into the equation

OpenStudy (freckles):

I'm not sure what you are picking... if you are trying to pick anything... 11=11(0) 12=11(1) 22=11(2) 1=11(11) --- the first equation is not true because 11(0)=0 and not 11 do you think the second equation is true? do you think 11 times 1 is 12?

OpenStudy (anonymous):

no its 11

OpenStudy (freckles):

right so the second equation is not true

OpenStudy (freckles):

how about the last two equations

OpenStudy (freckles):

is 22=11(2) is 1=11(11)

OpenStudy (anonymous):

As we all can see C is the only one that works

OpenStudy (anonymous):

because 121 does not equal 1

OpenStudy (anonymous):

is he right

OpenStudy (freckles):

yes 11 times 11 is 121 not 1

OpenStudy (wil476003):

It is equation c because if your put the numbers in for the variables the sides become equal

OpenStudy (freckles):

so the only true equation is 22=11(2)

OpenStudy (anonymous):

ok thanks

OpenStudy (freckles):

y=11x where y=22 and x=2 in other words y=11x is satisfied by (2,22)

OpenStudy (anonymous):

so its C

OpenStudy (freckles):

ye

OpenStudy (anonymous):

thx